40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 13 11:32:50 2008, Yoixhi, Update, PSL, MOPA is in trouble now powerdrop.png
    Reply  Mon Oct 13 13:56:03 2008, Yoichi, Update, PSL, MOPA is not that much in trouble now 
Message ID: 1044     Entry time: Mon Oct 13 13:56:03 2008     In reply to: 1042
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: MOPA is not that much in trouble now 
The problem was found to be all to do with the shutter.
The shutter started to work again, after a while, apparently for no clear reason.
The alignment to the PA was actually not screwed, and the MOPA output is now slowly increasing.
We figured that the 126MON PD has been mis-aligned for a long time. It was just picking the
scattered light from the output of the PA. So when the shutter is closed, it is natural that 126MON also goes down to zero.
It is a bit difficult to center the beam on the PD because there is not much room for moving the PD.
However, Alberto came up with a configuration (flip the PD and reflect back the beam with a mirror to the PD), which seems to
be feasible. We will do this modification when the MOPA is confirmed to be ok.

Here is more detail about the shutter problem:
The shutter is controlled by the MOPA power supply. There are three ways to command the power supply.
The switch on the front panel of the power supply, the EPICS switch (through a XYCOM XY220), and the interlock.
The ribbon cable from the power supply's back is connected to J1 of the cross connect. The pin 59 of the cable is the one
controlling the shutter. It is then routed to J12 pin 36. The interlock and a XYCOM switch are both connected to this
pin.
Now what happened was, on the way tracking down those cables, I pushed some connectors, especially the ones on the XYCOM.
After that, I was able to open the shutter from the EPICS button.
Steve and Alberto tried the EPICS button many times in the morning without success.
My guess is that it was some malfunctioning of the XY220 accidentally fixed by my pushing of the cables.
But I cannot exclude the possibility of the interlock malfunctioning.



Quote:
Steve, Alberto, Yoichi

A quick update.
The MOPA output went down to zero on Sunday early morning (00:28 AM).
We found that the NPRO beam is mis-aligned on the power monitoring PD (126MON).
We don't know yet if it is also mis-aligned to the power amplifier (PA) because the mechanical shutter is not working (always closed).
Most likely the beam is not aligned to the PA.
A mystery is that although the beam is terribly (more than a half inch) missing the monitor PD, the beam still goes through two faradays.
Another mystery is that the NPRO output power is now increased to 600mW.

The power drop was a very fast phenomenon (less than 1/16 sec).
We are trying to figure out what happened.
The first step is to fix the mechanical shutter. We have a spare in our hand.
ELOG V3.1.3-